OpenStudy (anonymous):

Whaaaaaatttt?? Pizza Planet delivers pizza within a 17 mile radius of their store. If this area is represented graphically, with Pizza Planet located at (2, 4), what is the equation that represents the delivery area?

OpenStudy (anonymous):

um... what is the scale? is every "one" on the graph one mile?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

well, lets assume that it is a one to one scale

OpenStudy (anonymous):

shall we?

OpenStudy (anonymous):

My answer options: A. (x + 2)^2 + (y + 4)^2 = 17 B. (x – 4)^2 + (y – 2)^2 = 117 C. (x – 2)^2 + (y – 4)^2 = 289 D. (x + 4)^2 + (y + 2)^2 = 289

OpenStudy (anonymous):

whaaaa?

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

If i had to guess, i would guess A, but I am not sure, ask someone else